# A Comprehensive Study on Large Language Models for Mutation Testing

### v3

@@ 3.3. RQ3: Performance on Efficiency
Average Generation Time: The row A. G. Time of Table 4 presents the results of Average Generation Time. Among the traditional appraoches, PIT demonstrates the fastest generation time at 0.02 seconds per mutation, followed by Major, which takes 0.08 seconds, making them the most efficient approaches. In contrast, LEAM and Bert take 3.06 seconds and 2.34 seconds, respectively. Among the LLMs, the LLMs of OpenAI (i.e., GPT-3.5, GPT-4o, and GPT-4o-Mini) show relatively faster generation, ranging from 1.31 to 1.79 seconds. While the open-source LLMs StarChat, CodeLlama, and DeepSeek require 7.53, 9.06, and 4.25 seconds, respectively. The results suggest that the rule-based approaches are the most time-efficient for generating mutations, which are tens to hundreds of times faster than other approaches.
@@ 3.4.1. Categorizations of compilation error types
Incorrect Method Parameters: this type of error is due to invoking a method with illegal parameters, including invoking by wrong numbers of parameters or unmatched type.
Usage of Unknown Variables: this type of error is due to using unknown variables under the code context.
Usage of Unknown Types: this type of error is due to using unknown type names under the code context.
@@
Incorrect Initialization: this type of error is due to using illegal forms of constrictors or deleting necessary initialization statements.
@@ 3.4.2. Categorizations of code context of non-compilable mutations
To analyze which types of code are prone to causing LLMs to generate non-compilable mutations, we examined the code locations of all non-compilable mutations generated by all nine mutation generation approaches in Table 4, as shown in Figure 3. For most approaches (except Major), the code locations with MemberReference and MethodInvocation are the top-2 prevent AST node types. For example, for GPT LLMs and DeepSeek, there about 40% of non-compilable mutations occur at the location with MemberReference or MethodInvocation. This is potentially caused by the inherent complexity of these operations, which often involve multiple dependencies and references. If any required method or member is missing or incorrectly specified, it can easily lead to non-compilable mutations. The errors highlight a need for better context-aware mutation generation, ensuring that method calls and member references align with the intended program structure. Additionally, we inspect the deletion mutations rejected by the compiler and find that they account for 7.48% (GPT-3.5), 1.69% (GPT-4o), 0.41% (GPT-4o-Mini), 1.11% (DeepSeek), 0.02% (StarChat), 0.14% (CodeLlama), 39.99% (LEAM), 0.11% (Bert), and 16.61% (Major) of all their non-compilable mutations, respectively. Thus, for LLMs, deletion is not the major reason for non-compilation. Thus for LLMs, deletion is not the major reason for non-compilation.
@@ 4.1.1. Setup of the Prompts
For the prompts exploration, we use two representative LLMs, i.e., the closed-source GPT-3.5-Turbo and the open-source DeepSeek-Coder-V2-236b, then switch different prompt templates. More specifically, we modify the default prompt template by adding and removing different sources of information in the Context, and we finally get four types of prompts, shown as follows:
@@
Prompt 4 (P4): P1 with Unit Tests. Based on the default prompt P1, P4 further adds the source code of corresponding unit tests of the target method into Context.
| Metric | GPT-3.5 | DS-236b | ||||||
| P1 | P2 | P3 | P4 | P1 | P2 | P3 | P4 | |
| Mut. # | 3681 | 3688 | 3350 | 3210 | 3803 | 3845 | 3802 | 3804 |
| Mut. Score | 0.773 | 0.785 | 0.842 | 0.801 | 0.784 | 0.778 | 0.793 | 0.814 |
| Real Bug Detec. | 79.4% | 78.4% | 72.2% | 72.5% | 84.6% | 84.0% | 81.4% | 82.6% |
| Avg. Ochiai | 63.1% | 61.8% | 60.8% | 62.1% | 67.3% | 64.1% | 65.0% | 63.7% |
| Coupling Rate | 52.1% | 52.2% | 47.4% | 49.6% | 53.1% | 51.9% | 52.9% | 50.7% |
| Compilability Rate | 66.9% | 66.8% | 60.0% | 61.2% | 79.1% | 78.9% | 74.9% | 78.2% |
| Dup. Mut. Rate | 11.4% | 13.3% | 14.9% | 12.7% | 10.9% | 9.4% | 9.8% | 12.7% |
| Avg. Gen. Time | 1.789 | 1.782 | 1.809 | 1.884 | 4.252 | 4.254 | 4.199 | 4.397 |
Table 12 presents a comparative evaluation of different prompts for GPT-3.5 and DeepSeek across multiple metrics. Among the four prompts, P1 consistently demonstrates superior performance across both models, achieving the highest Compilability Rates (66.9% for GPT-3.5 and 79.1% for DeepSeek) and the best Real Bug Detectability Rates (79.4% and 84.6%, respectively). P1 also yields the highest Avgerage Ochiai Coefficient across both models. These results overall indicate that providing LLMs with well-structured and appropriate information significantly enhances their performance (e.g., P1), whereas excessive information can lead to a decline in effectiveness (e.g., P4).
We further explore the performance similarity of different prompts. To measure their similarity in performance, we use the Spearman coefficient and Pearson correlation, as shown in the top part of Table 14. We can see that their mutual similarity exceeds 0.95, which is above the typical threshold of 0.85 (Siami Namin et al., 2008; Motwani and Brun, 2023; Kaufman et al., 2022), and the -values are all below 0.05.
@@ 4.1.2. Experiment Setup of the Context Length
| Settings | Spea. Coef. | -value | Pear. Coef. | -value |
@@
| GPT-3.5 P1 v.s. GPT-3.5 P4 | 0.9520 | 2.60E-04 | 0.9999 | 1.45E-25 |
| DS-236b P1 v.s. DS-236b P2 | 1.0000 | 0.00E+00 | 0.9999 | 1.94E-31 |
| DS-236b P1 v.s. DS-236b P3 | 0.9999 | 3.31E-05 | 0.9999 | 3.71E-30 |
| DS-236b P1 v.s. DS-236b P4 | 0.9762 | 3.31E-05 | 0.9999 | 1.83E-27 |
| 1-Line v.s. 2-Line | 0.9762 | 3.31E-05 | 0.9980 | 2.12E-08 |
| 2-Line v.s. 3-Line | 0.9762 | 3.31E-05 | 0.9997 | 1.04E-10 |
| 1-Line v.s. 3-Line | 0.9524 | 2.60E-04 | 0.9978 | 2.69E-08 |
| QB-6a (Default) v.s. QB-3 | 0.9701 | 6.55E-05 | 0.9996 | 1.70E-10 |
| QB-6a (Default) v.s. QB-3b | 0.9762 | 3.31E-05 | 0.9984 | 1.00E-08 |
| QB-6a (Default) v.s QB-9 | 0.9762 | 3.31E-05 | 0.9985 | 7.98E-09 |
@@
QB-N: Select N Examples from QuixBug. QB-6a is the set of default examples, while QB-6b involves 6 other examples. D4J-6 involves random 6 examples from Defects4J.
To compare the similarity to real defects, we deliberately select the context around the bug locations for generating mutations for all the approaches. To check the impact of context selection on performance, we conduct experiments on 60 bugs from Defects4J, comparing how different context lengths affect performance. In our previous experiments, the context length is set to three lines around the bug location, so we conduct experiments to compare the effects of two-line and one-line contexts, as shown in Table 13. Similarly, the middle part of Table 14 shows the performance similarity among different contexts. We can see that their mutual similarity exceeds 0.95 and the -values are all below 0.05. Different context lengths perform similarly, thus validating the setup of our experiment.
@@ 4.1.3. Impact of Different Few-Shot Examples
| Metric | QB-6a (Default) | QB-3 | QB-6b | QB-9 | D4J-6 |
| Real Bug Det. | 0.917 | 0.933 | 0.917 | 0.917 | 0.900 |
| Avg. Ochiai | 0.467 | 0.450 | 0.434 | 0.434 | 0.367 |
| Coupling Rate | 0.136 | 0.128 | 0.132 | 0.121 | 0.139 |
| Comp. Rate | 65.7% | 62.7% | 63.9% | 61.8% | 62.1% |
| Dup Mut. Rate | 13.7% | 12.8% | 12.8% | 13.8% | 12.5% |
| Equivalent Mutation Rate | 1.3% | 1.6% | 1.3% | 1.1% | 1.6% |
QB-N: Select N Examples from QuixBug. QB-6a is the set of default examples, while QB-6b involves 6 other examples. D4J-6 involves random 6 examples from Defects4J.
To measure the impact of different few-shot examples on the results, we conduct comparative experiments with various examples. First, we randomly select 3, 6, and 9 additional examples from QuixBugs, and ensure there is no overlap with the default 6 examples. Additionally, despite the risk of data leakage, we select 6 examples from the Defects4J dataset. The results are presented in Table 15.
We further compute their similarity, shown as the bottom part of Table 14. We can find that their similarity exceeds 0.95, significantly exceeding the typical threshold of 0.85, and the -values are all below 0.05. Therefore, we can conclude that although different few-shot examples have perturbation, their performance is very similar.

### v4

@@ 3.3. RQ3: Performance on Efficiency
Average Generation Time: The column Time of Table 4 presents the results of Average Generation Time in seconds. Among the traditional appraoches, PIT demonstrates the fastest generation time at 0.02 seconds per mutation, followed by Major, which takes 0.08 seconds, making them the most efficient approaches. In contrast, LEAM and BERT take 3.06 seconds and 2.34 seconds, respectively. The existing LLM-based approaches, Among the LLMs, the LLMs of OpenAI (i.e., GPT-3.5, GPT-4o, and GPT-4o-Mini) show relatively faster generation, ranging from 1.31 to 1.79 seconds. The open-source LLMs StarChat, CodeLlama, and DeepSeek require 7.53, 9.06, and 4.25 seconds, respectively. The results suggest that the rule-based approaches are the most time-efficient for generating mutations, which are tens to hundreds of times faster than other approaches. Note that the time costs of self-deployed open-source models typically depend on the types of GPUs used for execution. Additionally, the response time of models accessed via network APIs may vary across service providers and can be further affected by network fluctuations. We further discuss these potential impacts in the Section 5.
| Approach | Avg. Input Token | Avg. Output Token |
| BugFarm (GPT-4o) | 200 | 368 |
| LLMorpheus (GPT-4o) | 370 | 74 |
| LLMut (GPT-3.5) | 126 | 59 |
| LLMut (GPT-4o) | 126 | 59 |
| LLMut (GPT-4o-M) | 126 | 59 |
| LLMut (DS-761b) | 126 | 58 |
| LLMut (DS-236b) | 126 | 59 |
| LLMut (SC-16b) | 126 | 63 |
| LLMut (CL-13b) | 126 | 58 |
Average Token: Table 11 presents the average number of input and output tokens used by each LLM-based approach to generate one mutation. The existing LLM-based approaches, BugFarm and LLMorpheus, consume 200 and 370 input tokens, and 368 and 74 output tokens per mutation, respectively. LLMs using our prompt consume substantially fewer tokens, averaging 126 input tokens and 59–63 output tokens per mutation. As all LLMs share identical inputs, their input token counts remain the same.
@@ 3.4.1. Categorizations of compilation error types
Incorrect Method Parameters: This type of error is due to invoking a method with illegal parameters, including invoking with the wrong number of parameters or mismatched types.
Usage of Unknown Variables: This type of error is due to using unknown variables in the code context.
Usage of Unknown Types: This type of error is due to using unknown type names in the code context.
@@
Incorrect Initialization: This type of error is due to using illegal forms of constructors or deleting necessary initialization statements.
@@ 3.4.2. Categorizations of code context of non-compilable mutations
To analyze which types of code are prone to causing LLMs to generate non-compilable mutations, we examined the code locations of all non-compilable mutations generated by all nine mutation generation approaches in Table 4, as shown in Figure 3. For most approaches (except Major and BugFarm), the code locations with MemberReference and MethodInvocation are the top-2 most-prevent AST node types. For example, for GPT and DeepSeek models, about 40% of non-compilable mutations occur at the location with MemberReference or MethodInvocation. This is potentially caused by the inherent complexity of these operations, which often involve multiple dependencies and references. If any required method or member is missing or incorrectly specified, it can easily lead to non-compileable mutations. The errors highlight a need for better context-aware mutation generation, ensuring that method calls and member references align with the intended program structure. Additionally, we inspect the deletion mutations rejected by the compiler and find that they account for 7.48% (GPT-3.5), 1.69% (GPT-4o), 0.41% (GPT-4o-Mini), 3.4% (DeepSeek-V3-761b), 1.11% (DeepSeek-Coder-V2-236b), 3.4% (BugFarm), 0.8% (LLMorpheus), 0.02% (StarChat), 0.14% (CodeLlama), 39.99% (LEAM), 0.11% (BERT), and 16.61% (Major) of all their non-compilable mutations, respectively. Thus, for LLMs, deletion is not the major reason for non-compilation.
3.5. RQ5: Distributions of Alive Mutations
As shown in Table 4, the mutation scores of rule-based approaches (i.e., PIT and Major) are lower than LLMs, indicating their mutations are harder to detect. Therefore, we intend to analyze the reasons for survival mutations. To address this RQ, we adopt a setting similar to the analysis of equivalent mutations. Specifically, for each approach, we randomly sample its alive mutations and manually classify them. We employ the categories of alive mutation causes defined in the study by Just et al. (Just et al., 2014a) as follows.
(1) Not covered, i.e., the mutation is located in code never executed by the test suite. We refer to the mutation set of this category as .
(2) Covered but no state infection, i.e., the mutation is executed but does not propagate any state change. We refer to the mutation set of this category as .
(3) State infection without detection, i.e., the mutation alters the program state but any test oracle does not observe the deviation. We refer to the mutation set of this category as .
These categories are mutually exclusive, with the following relationships: , , and . This classification enables us to analyze the underlying reasons for alive mutations and further evaluate the extent to which different approaches contribute to strengthening the test suite. Given an alive mutation, we first determine whether it is equivalent; if so, it is discarded. For non-equivalent mutations, we then check whether they are uncovered. If a mutation is covered, we further examine whether it alters the program state, i.e., whether any memory values are modified. Finally, the remaining alive mutations that are covered but do not lead to detectable differences are classified as not detected. This stepwise process ensures that every alive mutation is uniquely assigned to one category (equivalent, uncovered, not injecting state, or not detected).
| Approach | #Sampled | #Eq. Mut. | NC | NI | ND | NC% | NI% | ND% |
| PIT | 96 | 3 | 43 | 27 | 23 | 46.24% | 29.03% | 24.73% |
| Major | 96 | 4 | 42 | 30 | 20 | 45.65% | 32.61% | 21.74% |
| LEAM | 95 | 7 | 59 | 16 | 13 | 67.05% | 18.18% | 14.77% |
| BERT | 88 | 8 | 37 | 24 | 19 | 46.25% | 30.00% | 23.75% |
| BugFarm (GPT-4o) | 94 | 9 | 62 | 11 | 12 | 72.94% | 12.94% | 14.12% |
| LLMopheus (GPT-4o) | 95 | 11 | 26 | 44 | 14 | 30.95% | 52.38% | 16.67% |
| LLMut (GPT-3.5) | 95 | 12 | 66 | 12 | 5 | 79.52% | 14.46% | 6.02% |
| LLMut (GPT-4o) | 95 | 11 | 53 | 16 | 15 | 63.10% | 19.05% | 17.86% |
| LLMut (GPT-4o-M) | 95 | 12 | 51 | 15 | 17 | 61.45% | 18.07% | 20.48% |
| LLMut (DS-761b) | 95 | 12 | 45 | 16 | 22 | 54.22% | 19.28% | 26.51% |
| LLMut (DS-236b) | 95 | 13 | 54 | 15 | 13 | 65.85% | 18.29% | 15.85% |
| LLMut (SC-16b) | 93 | 16 | 62 | 11 | 4 | 80.52% | 14.29% | 5.19% |
| LLMut (CL-13b) | 96 | 11 | 68 | 12 | 5 | 80.00% | 14.12% | 5.88% |
As shown in Table 13, the distributions of alive mutations differ significantly across approaches, where the columns NC%, NI%, and ND% represent the proportions of , , and within all non-equivalent mutations, respectively. First, LLM-based methods exhibit a substantially higher proportion of (i.e., not covered) mutations, meaning that more of their alive mutations reside in previously uncovered code lines. For instance, among the alive mutations generated by GPT-3.5 and BugFarm, 79.52% and 72.94% fall into the category, respectively, whereas PIT and Major exhibit much lower proportions of 46.24% and 45.65%. This characteristic highlights the potential of LLMs in guiding test suite augmentation by directing attention to untested program areas. Second, traditional approaches, especially rule-based methods such as PIT and Major, produce a larger fraction of (i.e., not injected) and (i.e., not detected) mutations compared to LLM-based methods. For instance, PIT and Major exhibit proportions of 29.03% and 32.61%, respectively, whereas all LLMs based on our prompt remain below 20%. These types of alive mutations require not only generating inputs that execute the mutated code but also propagating and detecting subtle semantic differences, which is inherently more challenging than increasing coverage. Therefore, the overall implication is that LLM-generated alive mutations are more directly beneficial for strengthening unit tests, as they emphasize extending coverage, while rule-based approaches tend to stress the harder task of state propagation and detection.
@@ 4.1.1. Setup of the Prompts
For the prompts exploration, we use two representative LLMs, i.e., the closed-source GPT-3.5-Turbo and the recent open-source model DeepSeek-V3-761b, then switch different prompt templates. More specifically, we modify the default prompt template by adding and removing different sources of information in the Context, and we finally get five types of prompts, shown as follows:
@@
Prompt 4 (P4): P1 with Unit Tests. Based on the default prompt P1, P4 further adds the source code of the corresponding unit tests of the target method into Context.
Prompt 5 (P5): P1 with Whole Java File. P5 is a variant of P1, which uses the code of the entire Java file as the code context, rather than the entire method.
| Metric | LLMut (GPT-3.5) | LLMut (DS-761b) | ||||||||
| P1 | P2 | P3 | P4 | P5 | P1 | P2 | P3 | P4 | P5 | |
| Mut. # | 3681 | 3688 | 3350 | 3210 | 3496 | 3851 | 3847 | 3834 | 3844 | 3878 |
| Mut. Score | 0.773 | 0.785 | 0.842 | 0.801 | 0.694 | 0.765 | 0.750 | 0.760 | 0.753 | 0.776 |
| Real Bug Detec. | 79.5% | 78.4% | 72.2% | 72.5% | 79.5% | 85.5% | 85.8% | 83.9% | 84.7% | 83.1% |
| Avg. Ochiai | 63.1% | 61.8% | 60.8% | 62.1% | 62.3% | 67.5% | 65.6% | 66.4% | 65.8% | 66.4% |
| Coupling Rate | 52.1% | 52.2% | 47.4% | 49.6% | 47.3% | 53.2% | 51.7% | 50.9% | 50.0% | 49.1% |
| Compling Rate | 66.9% | 66.8% | 60.0% | 61.2% | 70.3% | 82.8% | 83.5% | 78.0% | 82.0% | 79.5% |
| Dup. Mut. Rate | 11.4% | 13.3% | 14.9% | 12.7% | 16.9% | 7.3% | 6.4% | 8.5% | 7.0% | 8.2% |
| Avg. Gen. Time | 1.789 | 1.782 | 1.09 | 1.884 | 1.907 | 2.573 | 2.571 | 2.488 | 2.601 | 2.596 |
| Avg. # Token | 105 | 95 | 44 | 219 | 393 | 104 | 92 | 43 | 219 | 392 |
Note: The prompt template that achieves the best performance is highlighted in grey background color.
Table 14 presents a comparative evaluation of different prompts for GPT-3.5 and DeepSeek-V3-761b across multiple metrics. Across both models, P1 delivers the best overall performance on all effectiveness metrics (i.e., real bug detectability, average Ochiai, and coupling rate), demonstrating its strength in generating mutations that are semantically closer to real bugs. For the validity metrics, P1 ranks second in compilation rate on GPT-3.5 and first on DeepSeek-V3-761b, while achieving the lowest duplicate mutation rate on GPT-3.5 and third-lowest on DeepSeek-V3-761b. Notably, compared with P5, which provides the entire file as context, P1 relies only on the full method yet achieves higher semantic similarity to real bugs, with reduced generation time and token usage, while maintaining comparable compilation and duplicate mutation rates.
Across P1 to P5, we observe a clear context–performance trade-off. Removing relevant context (P3) substantially degrades effectiveness, as seen in lower real bug detectability, average Ochiai, and coupling rates. Providing a focused, method-level context (P1 and P2) delivers the best overall balance of effectiveness and validity. However, adding excessive context beyond the focal method (P4 and P5) does not yield further improvements and often increases token usage and generation time, while in some cases even reducing key effectiveness metrics. This suggests that moderate, relevant context benefits LLM mutation generation, whereas both insufficient and excessive context hurt efficiency and quality.
We further explore the performance similarity of different prompts. To measure their similarity in performance, we use the Spearman coefficient and Pearson correlation, as shown in the top part of Table 16. We can see that their mutual similarity exceeds 0.95, which is above the typical threshold of 0.85 (Siami Namin et al., 2008; Motwani and Brun, 2023; Kaufman et al., 2022), and the -values are all below 0.05. Although P3 achieves the most balanced overall performance, the differences among prompts are not statistically significant.
@@ 4.1.2. Experiment Setup of the Context Length
| Settings | Spea. Coef | -value | Pear. Coef | -value |
@@
| GPT-3.5 P1 v.s. GPT-3.5 P4 | 0.9520 | 2.60E-04 | 0.9999 | 1.94E-31 |
| GPT-3.5 P1 v.s. GPT-3.5 P5 | 0.9762 | 3.31E-05 | 0.9999 | 1.57E-26 |
| DS-671b P1 v.s. DS-671b P2 | 1.0000 | 0.00E+00 | 0.9999 | 1.08E-31 |
| DS-671b P1 v.s. DS-671b P3 | 1.0000 | 0.00E+00 | 0.9999 | 2.93E-29 |
| DS-671b P1 v.s. DS-671b P4 | 1.0000 | 0.00E+00 | 0.9999 | 4.95E-31 |
| DS-671b P1 v.s. DS-671b P5 | 1.0000 | 0.00E+00 | 0.9999 | 1.51E-30 |
| 1-Line vs. 2-Line | 0.9762 | 3.31E-05 | 0.9980 | 2.12E-08 |
| 2-Line vs. 3-Line | 0.9762 | 3.31E-05 | 0.9997 | 1.04E-10 |
| 1-Line vs. 3-Line | 0.9524 | 2.60E-04 | 0.9978 | 2.69E-08 |
| QB-6a (Default) v.s. QB-6b | 0.9429 | 4.81E-03 | 0.9984 | 1.00E-08 |
| QB-6a (Default) v.s. QB-3 | 0.9856 | 3.09E-05 | 0.9999 | 1.62E-13 |
| QB-6a (Default) v.s. QB-4 | 0.9643 | 4.54E-04 | 0.9999 | 7.28E-23 |
| QB-6a (Default) v.s. QB-5 | 0.9643 | 4.54E-04 | 0.9999 | 9.62E-24 |
| QB-6a (Default) v.s. QB-7 | 1.0000 | 0.0E+00 | 0.9999 | 9.39E-16 |
| QB-6a (Default) v.s. QB-8 | 1.0000 | 0.0E+00 | 0.9999 | 3.34E-15 |
| QB-6a (Default) v.s. QB-9 | 1.0000 | 0.00E+00 | 0.9999 | 5.01E-14 |
| QB-6a (Default) v.s. QB-10 | 0.9856 | 3.09E-04 | 0.9999 | 2.89E-13 |
@@
QB-N: Randomly samples N Examples from QuixBug. QB-6a is the set of default examples, while QB-6b randomly samples 6 other examples. D4J-6 randomly samples 6 examples from Defects4J.
To compare the similarity to real defects, we deliberately select the context around the bug locations for generating mutations for all the approaches. To check the impact of context selection on performance, we conduct experiments on 60 bugs from Defects4J, comparing how different context lengths affect performance. In our previous experiments, the context length is set to three lines around the bug location, so we conduct experiments to compare the effects of two-line and one-line contexts, as shown in Table 15. Similarly, the middle part of Table 16 shows the performance similarity among different contexts. We can see that their mutual similarity exceeds 0.95, and the -values are all below 0.05. Different context lengths perform similarly, thus validating the setup of our experiment.
@@ 4.1.3. Impact of Different Few-Shot Examples
| Metric | QB-6a | QB-3 | QB-4 | QB-5 | QB-6b | QB-7 | QB-8 | QB-9 | QB-10 | D4J-6 |
| Real Bug Detec | 0.917 | 0.933 | 0.883 | 0.883 | 0.917 | 0.900 | 0.900 | 0.917 | 0.867 | 0.900 |
| Avg. Ochiai | 0.467 | 0.450 | 0.400 | 0.433 | 0.433 | 0.467 | 0.433 | 0.433 | 0.400 | 0.367 |
| Coupling Rate | 0.136 | 0.128 | 0.126 | 0.127 | 0.132 | 0.112 | 0.113 | 0.121 | 0.136 | 0.139 |
| Comp. Rate | 67.5% | 62.7% | 65.6% | 66.8% | 63.9% | 65.1% | 66.0% | 61.8% | 64.9% | 62.1% |
| Dup.Mut.Rate | 13.7% | 12.8% | 12.6% | 9.4% | 12.8% | 11.4% | 13.2% | 13.8% | 13.6% | 13.9% |
| Avg. # Token | 106 | 98 | 102 | 103 | 106 | 108 | 109 | 110 | 113 | 107 |
QB-N: Randomly samples N Examples from QuixBug. QB-6a is the set of default examples, while QB-6b randomly samples 6 other examples. D4J-6 randomly samples 6 examples from Defects4J.
To measure the impact of different few-shot examples on the results, we conduct comparative experiments with various examples. First, we randomly select 3, 4, 5, 6, 7, 8, 9, and 10 additional examples from QuixBugs, and ensure there is no overlap with the default 6 examples. Additionally, despite the risk of data leakage, we select 6 examples from the Defects4J dataset. The results are presented in Table 17.
We further compute their similarity, shown as the bottom part of Table 16. We can find that their similarity exceeds 0.95, significantly exceeding the typical threshold of 0.85, and the -values are all below 0.05. Therefore, despite variations in few-shot examples, there is no statistically significant evidence indicating superior performance among the prompts.